How many solutions does the following system have:
6x - 3y = 9
2x - y = 3
Explain how to determine the number of solutions without solving the system. Then apply elimination, and interpret the resulting equation.

Answer by checkley79(3341)

X & Y should have one solution each.
However in this case - there are no solutions for x or y because these two eauations are the same.
6x - 3y = 9
2x - y = 3 MULTIPLY THIS EQUATION BY -3 & ADD.
6x-3y=9
-6x+3y=-9
-----------
0x+0y=0